OpenStudy (anonymous):

A rectangular prism has a volume of 260 in.³ Its width is 5 in. and its length is 8 in. The height of the prism is _____ in.

OpenStudy (anonymous):

v=lwh

OpenStudy (anonymous):

260=8*5*x 260=40x 260/40=x